### v3

# TIMeSynC: Temporal Intent Modelling with Synchronized Context Encodings for Financial Service Applications

Abstract.
Users engage with financial services companies through multiple channels, often interacting with mobile applications, web platforms, call centers, and physical locations to service their accounts. The resulting interactions are recorded at heterogeneous temporal resolutions across these domains. This multi-channel data can be combined and encoded to create a comprehensive representation of the customer’s journey for accurate intent prediction. This demands sequential learning solutions. NMT transformers achieve state-of-the-art sequential representation learning by encoding context and decoding for the next best action to represent long-range dependencies. However, three major challenges exist while combining multi-domain sequences within an encoder-decoder transformers architecture for intent prediction applications: a) aligning sequences with different sampling rates b) learning temporal dynamics across multi-variate, multi-domain sequences c) combining dynamic and static sequences. We propose an encoder-decoder transformer model to address these challenges for contextual and sequential intent prediction in financial servicing applications. Our experiments show significant improvement over the existing tabular method.
sequential recommendation, tabular synthesis, transformers, temporal learning, heterogeneous data
copyright: acmlicensedjournalyear: 2024conference: Eighteenth ACM Conference on Recommender Systems; Oct 14–18, 2024; Bari, Italyccs: Computing methodologies Machine learningccs: Information systems Recommender systemsccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Temporal data
1. Introduction
Companies in the financial services industry have adopted a multi-channel approach to enable customers to service their accounts including shopping at stores. User behavior across all channels can be used to deliver a personalized digital experience at the moment of sign-on as well as in-session on digital platforms for targeted marketing, increased find-ability, contextual chatbot Q&A, and more. In financial services, the data is stored in tabular format (Qiu et al., 2019), but it also exhibits strong sequential properties. Unlike tabular data where rows are independent, sequential tabular data exhibits dependent rows with dependent columns (Zhang et al., 2023; Luetto et al., 2023). This data exhibits strong dynamic and static patterns capturing local transient patterns and global identity (Padhi et al., 2021). For example, the customer’s click-stream activity and the amount spent exhibit a dynamic pattern, while the type of product owned exhibits a static pattern replicated across the sequence for an elongated period. Another property of multi-domain data is to be stored at the granular level of the respective timeline as illustrated in Figure 2. These properties make an important representation of a customer’s state for their next best action. It’s a common practice to perform feature engineering and convert raw sequences into the tabular data (Kotios et al., 2022; Borisov et al., 2023) for machine learning algorithms such as Gradient Boosted Trees (Kotios et al., 2022; Borisov et al., 2023; Taha and Malebary, 2020). However, extensive feature engineering is required on sequential tabular data leading to long hours spent in developing the accurate features. Similarly, methods for representing the customer’s next action based on the customer’s sequence of historical interactions are well explored (Kang and McAuley, 2018; Sun et al., 2019; Pancha et al., 2022; Hidasi and Karatzoglou, 2018; Cao and Lio, 2024; de Souza Pereira Moreira et al., 2021; Moreira et al., 2021; Yuan et al., 2020; Chang et al., 2021). For contextual sequential recommendation, transformers have often shown superior performance over other architectures (Kang and McAuley, 2018; Sun et al., 2019; de Souza Pereira Moreira et al., 2021; Chen et al., 2024). However, due to the cross-correlation between sequences and miss-alignment, out-of-the-box transformer architecture cannot be directly applied. Another challenge is to incorporate time information into the sequence to better understand customer’s seasonal, periodic, and temporal behavior across the domains.
2. Relevant Work
2.1. Sequential Recommendation
The evolution of sequential recommendation systems has been driven by advancements in deep learning, particularly in the ability to model user behavior over time. Sequential recommendation models initially focused on capturing patterns in user behavior using techniques like Markov Chains (MC) and Recurrent Neural Networks (RNN) but struggles with long-term dependencies and computational efficiency. SASRec (Kang and McAuley, 2018) introduced self-attentive attention to capturing both short-term actions and long-term user preferences, lacking a bi-directional representation of the events. BERT4Rec (Sun et al., 2019) extended SASRec by introducing masked loss and allowing the model to represent future actions. Transformer-based models showed success in some industries where recommendation system applications have high business value, most evident in the e-commerce sector, where user behavior data is abundant and continually evolving; for instance Alibaba’s Behavior Sequence Transformer (BST) (Chen et al., 2019), Pinterest’s TransAct (Xia et al., 2023; Pancha et al., 2022), LinkedIn’s LiGNN (Borisyuk et al., 2024) and many others. These innovations address practical challenges like scalability, latency, and cost, but do not do enough to address heterogeneous data within the contextual recommendation framework where heterogeneous data is abundant.
2.2. Handling Tabular Data with LLM
Recent advancements in large language models (LLMs) have opened new opportunities for handling structured data, notably in tabular data synthesis and understanding. TaBert (Yin et al., 2020) extends the capabilities of BERT to tabular data, by introducing a content snapshot mechanism, that efficiently encodes relevant table segments to align with natural language queries, significantly improving processing speed and accuracy. TabuLa (Zhao et al., 2023) introduces a novel synthesis framework for tabular data, by employing a specialized compression strategy and a unique padding method and addressing the efficiency and quality of synthetic data generation. Finally, the Table-GPT (Zha et al., 2023) framework represents a significant leap in the table-specific tuning of generative pre-trained transformers. Together, these models illustrate the transformative impact of LLMs in the field of tabular data handling. Even though they allow for more efficient table representation, to our knowledge there are no benchmarks available for representing heterogeneous data in financial services applications for sequential recommendation.
Figure 1. Model Architecture. Field Value embeddings are added with Product and Field Name embeddings before employing the encoder block. For both the decoder and encoder, a Time encoder and TimeAliBi are applied for learning absolute and relative time dynamics. Depending upon the self-attention or cross-attention, a causal mask(K(t), Q(t)) is applied with respect to the time corresponding to Key(K) and the Query(Q). In blue the encodings used as input of the Encoder block; in Red the TimeAliBi self-attention module; in purple the input encodings to the Decoder module; in yellow and green the decoder TimeAlibi cross-attention modules, and at last in white the head layers. This figure only notes changes w.r.t. to the vanilla transformers (Vaswani, 2017)
2.3. Sequential Tabular Learning and Representation
The importance of tabular data sequences as a common data storage type to represent financial transactions recorded in a bank database (Luetto et al., 2023), has led researchers to improve modeling (Arik and Pfister, 2021). TabNET (Arik and Pfister, 2021) introduced a deep learning architecture focused on tabular structures. It incorporates sequential attention outperforming both traditional machine learning and early deep learning. TabTransformer (Gorishniy et al., 2021) expands TabNET by integrating a Transformer Encoder block for domain-specific field embeddings, followed by FT-Transformer to contextualize continuous and numerical features. However, it did not sufficiently address the sequential aspect of the data. TabBERT (Padhi et al., 2021), adapted from the BERT architecture (Devlin, 2018), is a hierarchical architecture that utilizes self-attention mechanisms to capture complex relationships between features and temporal dependencies in sequential tabular data, demonstrating improved performance on financial forecasting and customer behavior prediction tasks. FATA-Trans (Zhang et al., 2023) further extends this approach by isolating static tabular features from dynamic sequential data, enabling better differentiation between static and dynamic fields, reducing negative effects from static field replication, and improving temporal information learning. Similarly, UnitTAB (Luetto et al., 2023) extends TabBERT by incorporating row-type dependent embedding to handle variable data dimensions. The model (Yin et al., 2020; Luetto et al., 2023; Zhang et al., 2023) are reported to be used for multiple classification tasks such as pollution prediction, fraud detection, and loan default prediction.
In this context, we propose a novel approach called TIMeSynC (Temporal Intent Modelling with Synchronized Context Encodings) for contextual recommendations in financial services that incorporates the following items:
  • First flatten and tokenize the customer’s multi-channel activity as an encoder context and incorporate a sequence of online intent as a decoder for the next intent prediction.
  • ALiBi (Press et al., 2021) based time representation and time encoder for efficient learning of the temporal cross-correlation between multivariate and multi-domain sequences.
  • A time-based attention mask for cross-attention in the decoder module is employed to dynamically attend causal tokens for learning temporal patterns between intent and the context.
  • Field Name embeddings are added to the encoder for domain and field-aware representation.
  • Static sequences of financial product enrollment are fused for point-in-time representation with the decoder output. Product embeddings are added to the encoder sequence for product-aware representation.
These additional components eliminate the need for multi-level hierarchical modeling, simplifying the architecture to learn complex, deeper, and richer patterns by direct interaction between cross-field and cross-domain dependencies over time.
Table 1. Sample intents
Make a payment Report Fraud Cancel Zelle Payment
Redeem rewards Add External Account
Enroll into paperless Apply Contactless Card
Inquire about Benefits Cancel Balance Transfer Update Birthday
Table 2. Sample data streams used for this study
Domains Description
Transactions Credit Card transactions
Payments Credit Card payments
Rewards Rewards redeemed and earned activities
Outbound messages Servicing messages to the customer
Product Enrollment Point in time customer product enrollment
3. Problem Statement
The sequential intent prediction problem in finance aims to forecast user actions based on inputs from diverse domains. Let $X=[X_{1},X_{2},\ldots,X_{n}]$ represent $n$ domains, where $X_{i}$ denotes different types of financial interactions, such as credit card/debit card transactions, payments, rewards, account status, customer care messages, call center intents, digital page views, etc. Each $X_{i}$ is a time series represented as $[X_{i1},X_{i2},\ldots,X_{it}]$. The goal is to predict future user’s intents $Y=[y_{1},y_{2},\ldots,y_{k}]$ at time $t$, given the financial context $X$ and the prior sequence of user’s intents $Y$.
Figure 2. Illustrative customer journey across domains at their respective timelines
Figure 3. Illustrative Example. Data Flattening and Tokenization of the Encoder Context
4. Experiments
4.1. Experimental Setup
We use various data sources from the internal data ecosystem as mentioned in Table 2. Each data source is associated with its respective timestamps. Therefore, we can construct a flat sequence of customer’s activities all arranged by their respective timestamp. An example of a customer timeline is shown in Figure 2. We then temporally split the intent and context sequence per user into training/validation/test for model training. We focus our research on 30M+ user sequences with at least one financial activity in the past, and using 500+ intents as supervision. Table 1 shows the sample intent space for this study.
Finally, Assuming we have a set of user U for each intent I, we compute for evaluation the top k recall as:
[…]
4.2. Model Framework
4.2.1. Baseline(s)
Table 3. Offline relative lift in ranking recall over SASRec expressed in %. Ranking recall is calculated by sorting softmax scores in descending order
Methods Recall@1 Recall@5 Recall@10
SASRec 0.00 0.00 0.00
SASRec + Tabular Context +2.60 +2.23 +0.03
SASRec + Encoder(Tabular Context) +4.23, +1.59 +4.06, +1.79 +0.98, +0.95
TIMeSynC +7.81, +3.43 +5.54, +1.43 +2.18, +1.19
  • SASRec (Kang and McAuley, 2018): a self-attention based causal recommender, that estimates the importance of the sequence itself without the context. We employ learned positional embeddings as described in (Kang and McAuley, 2018).
  • SASRec + Tabular Context: This is our current method. We perform extensive feature engineering across different data domains to represent customer’s financial context state over time in tabular format. These features are aligned with the timestamp of the intent. We then fuse these features with the SASRec output before the final loss.
  • SASRec + Encoder(Tabular Context): To measure existing gaps in our feature engineering, we add an encoder block for the tabular features. We then adapt causal encoder-decoder architecture from (Vaswani, 2017). For the encoder module, a causal time mask is applied based on the context timestamp. For the cross-attention, a causal time mask is applied based on context and intent timestamp.
Table 4. Relative negative effect of each feature when excluded from training expressed in %
Feature Recall@1
Full Model 100.00
Decoder w/o Time-stamp Encoder -0.227
Decoder w/o TimeAliBi Self Attn. -0.361
Decoder w/o TimeAliBi Cross Attn. -0.715
Decoder w/o Product Embedding -1.246
Encoder w/o Field Name Embedding -1.272
Encoder w/o Time-stamp Encoder -0.815
Encoder w/o TimeAliBi Self Attn. -1.082
Encoder w/o Product Embeddings -1.338
4.2.2. TIMeSynC:
For each data source, we first quantile-discretize each numerical field in bins and convert it to string. Then we combine and collapse different data sources together. Finally, we sort them by their respective timestamps as summarized in Figure 3. This results in X(u) with u the User and n the sequence length:
$[(t_{i},...,t_{n}),(fn_{i},...,fn_{n}),(fv_{i},...,fv_{n}),(p_{i},...,p_{n})]$
Where t is the timestamp, fn is the field name, fv is the field value, and p is the product. All fields in X are then string-tokenized except the timestamp. Similarly, sequences of intents are sorted by the timestamp are then string tokenized resulting in Y(u):
$[(t_{i},...,t_{n}),(y_{i},...,y_{n})]$
Where t is the timestamp, y is the intent. The model architecture is visualized in Figure 1 and described below:
  • TimeAliBi($t(m)$, $t(n)$): Inspired by AliBi (Press et al., 2021), we modify the attention mechanism (Vaswani, 2017) to include a bias based on relative time. Before computing the final attention scores, a relative time matrix is added to the attention scores, followed by applying the softmax function for score normalization. Formally, we define:
    $\text{softmax}\left(\frac{q_{i}K^{\mathrm{T}}+s\cdot[t(q_{i})-t(k_{j})\mid i\in\{1,\ldots,n\},j\in\{1,\ldots,m\}]}{\sqrt{d_{k}}}\right)\cdot V$
    where $V$ is the value matrix, $q_{i}$ is the $i$-th query, $K$ is the key matrix, $n$ is the query sequence length, $m$ is the key sequence length, and $s$ is the slope parameter.
  • Context Encoder: The Transformer Encoder Layer, as described in (Vaswani, 2017), processes a raw sequence of field values ($S$). A temporal causal mask is then applied based on relative time. To capture sequence periodicity, seasonality, and frequent behavior, time representations based on multi-dimensional signals such as the day of the week, week of the month, and hour of the day are added to the sequence embeddings as described in (Rahmani et al., 2023). The TimeAliBi($E(t)$, $E(t)$) function is then used to capture the relativity of time, where $E(t)$ denotes the time associated with the encoder sequence.
[…]
  • Field Name Embeddings: Due to flattening and tokenization, domain and field representations can be lost for certain quantities (e.g., amount). To retain this information, we encode field and domain names as <field_name_+_domain>. These are then string-tokenized and embedded before being added to the raw sequence.
  • Product Representation: Users can own multiple accounts, reflecting their interactions with those accounts. To incorporate product awareness in the encoder, product embeddings are added to the raw sequence of the encoder. Additionally, to identify point-in-time product ownership in the decoder module, the product one-hot encoding is first passed through a dense layer and then fused with the final decoder output.
4.3. Results
Table 4 shows a performance comparison between TIMeSynC and baselines. We specifically consider short-term ranking recall@1,5,10 due to the business needs and the UI specifications. The tabulated values are expressed as %, the best results are boldfaced and the relative lift from the prior row is underlined. SASRec provides an important baseline for understanding the value of a sequence of intents and establishing a lower bound. Both point-in-time and sequential context with the encoder demonstrate superior performance over SASRec showing the importance of tabular context. TIMeSynC outperforms SASRec and further shows improved performance over SASRec + Encoder(Tabular Context) proving the loss of signal in prior methods. Feature Ablation: In Table 4, we see the impact of each feature on the final model performance by removing it from training. The two features that contribute noticeably are the Product embedding and Field Name embedding. The temporal aspects are of higher importance in the encoder context than the decoder block. Given the importance of each feature, we choose to include all features in TIMeSynC.
5. Conclusion and Future Work
In this paper, we proposed a novel framework for context-aware sequential recommendation in financial services applications for effectively combining heterogeneous data with sequential actions. Specifically, we tackle this problem by leveraging an encoder-decoder architecture and flattening across the domain, field, and time. We also employ TimeAliBi and a multi-dimensional time encoder for representing absolute and relative time. Our empirical results demonstrate the gap in the existing feature-engineered tabular context and highlight the significance of our approach. Limitation(s): We acknowledge that our approach for encoding context has some constraints compared to a hierarchical framework approach, due to the flattening across the domain, field, and time that could lead to an explosion of the encoder context window. Additionally, we acknowledge that the results may not generalize to other datasets that do not exhibit the same characteristics. In the Future: We aim to improve numerical representation, and tokenization and incorporate other data sources such as click-stream, call center, and credit bureau data. We further aim to apply TIMeSynC to other recommendation objectives (vehicle and marketing recsys, mobile app, shopping, and business deals personalization). Broader Impact: While our approach is primarily studied in the context of user’s Q&A intent prediction, it may apply to other prominent objectives in financial services such as fraud, call center intent, and charge-off prediction. Our work could apply to other industries like e-commerce, entertainment, and tourism given the peculiarities and differences in data.
Acknowledgements.
We would like to thank Arturo Hernandez Zeledon who contributed to the discussions and reviewed the paper. We would like to thank Qi Yu and Giri Iyengar who supported us through this study.